• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Nuruzzaman

    Log entry time 15:57:48 on September19,2009

    Entry number 290538

    keyword=Day Shift Summary

    08.10 Shift started with production data @ 98uA, LH2, HW plate out.
    09.10 Compton CODA restarted several times,finally worked.
    09.45 Started compton run 19514 with config. "FADCScaler".
    11.48 Happex CODA lost communication with ROC31. 
    11.57 Changed config. to HapNoQweak from "happex".
    15.49 Started parity run 12492
    ------------------------------------------------
    Parity Rns(production data)			
    Time	Run#	Beam(uA) Total_events	OK_pairs
    ------------------------------------------------
    07.56	12481	 98  	  100k		  43.3k
    08.53	12482	 98 	  100k		  37.9k
    09.50	12483	 98	  100k		  44.5k
    10.47	12484	 98	  100k		  42.7k
    11.59	12488	 98	  100k		  43.5k
    12.56	12489	 98	  100k		  44.9k
    13.53	12490	 98	  100k		  43.8k
    14.52	12491	 98	  100k		  44.8k
    ------------------------------------------------
    sopt++ run
    Time	Run#  Total_events Raster
    ------------------------------------------------
    14.50	24767	53k	   6.5x5.0
    
    
    Comments: Mysterious HV alarms few times during the shift.
    	  Target behaved well.
    	  RMS lumi asymmetries were around 800ppm, expected 1100ppm
    	  Quite shift with good production data.